# Text-Trained LLMs Can Zero-Shot Extrapolate PDE Dynamics

### v1

@@ Abstract
Large language models (LLMs) have demonstrated emergent in-context learning (ICL) capabilities across a range of tasks, including zero-shot time-series forecasting. We show that text-trained foundation models can accurately extrapolate spatiotemporal dynamics from discretized partial differential equation (PDE) solutions without fine-tuning or natural language prompting. Predictive accuracy improves with longer temporal contexts but degrades at finer spatial discretizations. In multi-step rollouts, where the model recursively predicts future spatial states over multiple time steps, errors grow algebraically with the time horizon, reminiscent of global error accumulation in classical finite-difference solvers. We interpret these trends as in-context neural scaling laws, where prediction quality varies predictably with both context length and output length. To better understand how LLMs are able to internally process PDE solutions so as to accurately roll them out, we analyze token-level output distributions and uncover a consistent ICL progression: beginning with syntactic pattern imitation, transitioning through an exploratory high-entropy phase, and culminating in confident, numerically grounded predictions.
1 Introduction
@@
Large language models (LLMs) exhibit an emergent ability known as in-context learning (ICL) (Brown et al., 2020; Dong et al., 2022; Zhao et al., 2025), in which the model is conditioned on a sequence of examples and/or task instructions provided in the input and learns to generate appropriate outputs for new instances—without any parameter updates or additional training on task-specific data. In the zero-shot setting, LLMs are given only a task description and/or a serialized input and are expected to generalize purely from the prompt.
@@
While ICL was initially observed in linguistic tasks, it has since been demonstrated in domains involving mathematical reasoning (Wei et al., 2022a; b; Akyürek et al., 2023; Garg et al., 2022). Recent work shows that LLMs such as GPT-3 (Brown et al., 2020) and Llama-2 (Touvron et al., 2023) can, in the zero-shot setting, forecast time series (Gruver et al., 2023; Jin et al., 2024), infer governing principles of dynamical systems (Liu et al., 2024), and perform regression and density estimation (Requeima et al., 2024; Liu et al., 2025). From a theoretical perspective, in-context scaling laws have been analyzed by modeling LLM inference as a finite-state Markov chain, yielding analytical results for Markov-chain–generated inputs (Zekri et al., 2024), and by developing theoretical explanations of ICL scaling behavior when LLMs learn Hidden Markov Models (Dai et al., 2025).
@@
We demonstrate that pretrained LLMs, such as Llama-3 (Grattafiori et al., 2024), Phi-4 (Abdin et al., 2024), and SmolLM3 (Hugging Face, 2025), possess an additional zero-shot ICL capability: the ability to continue the dynamics of partial differential equations (PDEs) directly from serialized solution data (see Section 3). Our focus is on time-dependent PDEs, whose solutions often exhibit multi-dimensional correlations, long-range dependencies, and stiff nonlinear dynamics (Evans, 2010; Haberman, 2013). We adopt the following setup: representing spatiotemporal data as delimited sequences of real numbers and feeding them directly into an LLM, without any fine-tuning or natural-language prompting. The model generates token sequences autoregressively, effectively learning to infer both spatial structure and temporal dynamics from in-context information alone (see Figure 1). We emphasize that we do not propose to employ LLMs as a new kind of PDE solver. Instead, we study their ICL behavior in continuing the spatiotemporal dynamics of PDEs, as a lens to investigate the inductive biases and numerical priors that emerge from large-scale pretraining. Python code to reproduce all numerical experiments is publicly available at https://github.com/Jiajun-Bao/LLM-PDE-Dynamics.
Main Contributions.
@@
2 Background
Recent work at the intersection of LLMs and PDEs follows two main directions: (i) using LLMs as assistants in scientific modeling pipelines, and (ii) leveraging LLMs as direct PDE solvers. We briefly review representative examples from each line of research.
LLMs as Assistants in PDE Pipelines. Jiang et al. (2025) evaluate LLMs on tasks such as implementing numerical solvers and constructing scientific machine learning pipelines. Li et al. (2025) introduce CodePDE, a framework that formulates PDE solving as code generation. Soroco et al. (2025) propose PDE-Controller, which enables LLMs to convert informal natural language instructions into formal specifications for PDE control. Lorsung & Farimani (2024) leverage LLMs to integrate prior knowledge to improve PDE surrogate models. Zhou et al. (2025) present Unisolver, a neural PDE solver conditioned on symbolic PDE embeddings produced by LLMs.
LLMs as PDE Solvers. The Universal Physics Solver (UPS) (Shen et al., 2024) adapts pretrained LLMs to learn unified neural solvers for time-dependent PDEs. ICON-LM (Yang et al., 2025) fine-tunes LLMs for in-context operator learning. These methods typically involve custom architectures or task-specific training procedures tailored for PDE solving.
@@ 2 Background
3 Methodology

### v2

@@ Abstract
Large language models (LLMs) have demonstrated emergent in-context learning (ICL) capabilities across a range of tasks, including zero-shot time-series forecasting. We show that text-trained foundation models can accurately extrapolate spatiotemporal dynamics from discretized partial differential equation (PDE) solutions without fine-tuning or natural language prompting. Predictive accuracy improves with longer temporal contexts but degrades at finer spatial discretizations. In multi-step rollouts, where the model recursively predicts future spatial states over multiple time steps, errors grow algebraically with the time horizon, reminiscent of global error accumulation in classical finite-difference solvers. We interpret these trends as in-context neural scaling laws, where prediction quality varies predictably with both context length and output length. To better understand how LLMs are able to internally process PDE solutions so as to accurately roll them out, we analyze token-level output distributions and uncover a consistent three-stage ICL progression: beginning with syntactic pattern imitation, transitioning through an exploratory high-entropy phase, and culminating in confident, numerically grounded predictions.
keywords:
Large language models, in-context learning, partial differential equations, interpretability
Introduction
@@
Large language models (LLMs) exhibit an emergent ability known as in-context learning (ICL) [1, 2, 3], in which the model is conditioned on a sequence of examples and/or task instructions provided in the input and learns to generate appropriate outputs for new instances—without any parameter updates or additional training on task-specific data. In the zero-shot setting, LLMs are given only a task description and/or a serialized input and are expected to generalize purely from the prompt.
@@
While ICL was initially observed in linguistic tasks, it has since been demonstrated in domains involving mathematical reasoning [4, 5, 6, 7]. Recent work shows that LLMs such as GPT-3 [1] and Llama-2 [8] can, in the zero-shot setting, forecast time series [9, 10], infer governing principles of dynamical systems [11], and perform regression and density estimation [12, 13]. From a theoretical perspective, in-context scaling laws have been analyzed by modeling LLM inference as a finite-state Markov chain, yielding analytical results for Markov-chain–generated inputs [14], and by developing theoretical explanations of ICL scaling behavior when LLMs learn Hidden Markov Models [15].
@@
We demonstrate that pretrained LLMs, such as Llama-3 [16], Phi-4 [17], and SmolLM3 [18], possess an additional zero-shot ICL capability: the ability to continue the dynamics of partial differential equations (PDEs) directly from serialized solution data (see Methods section). Our focus is on time-dependent PDEs, whose solutions often exhibit multi-dimensional correlations, long-range dependencies, and stiff nonlinear dynamics [19, 20]. We adopt the following setup: representing spatiotemporal data as delimited sequences of real numbers and feeding them directly into an LLM, without any fine-tuning or natural-language prompting. The model generates token sequences autoregressively, effectively learning to infer both spatial structure and temporal dynamics from in-context information alone (see Fig. 1). We emphasize that we do not propose to employ LLMs as a new kind of PDE solver. Instead, we study their ICL behavior in continuing the spatiotemporal dynamics of PDEs, as a lens to investigate the inductive biases and numerical priors that emerge from large-scale pretraining. Our key contributions are as follows:
@@
Background
Recent work at the intersection of LLMs and PDEs mainly follows two directions: (i) using LLMs as assistants in scientific modeling pipelines, and (ii) employing LLMs as direct PDE solvers. We briefly review representative examples from each line of research.
LLMs as Assistants in PDE Pipelines.
Jiang et al.[21] evaluate LLMs on tasks such as implementing numerical solvers and constructing scientific machine learning pipelines. Li et al.[22] introduce CodePDE, a framework that formulates PDE solving as code generation. Soroco et al. [23] propose PDE-Controller, which enables LLMs to convert informal natural language instructions into formal specifications for PDE control. Lorsung & Farimani [24] leverage LLMs to integrate prior knowledge to improve PDE surrogate models. Zhou et al.[25] present Unisolver, a neural PDE solver conditioned on symbolic PDE embeddings produced by LLMs. Zhou et al.[26] develop Text2PDE, a diffusion–based neural PDE solver that supports text-conditioned simulation using captions generated by a multimodal LLM.
LLMs as PDE Solvers.
The Universal Physics Solver (UPS) [27] adapts pretrained LLMs to learn unified neural solvers for time-dependent PDEs. ICON-LM [28] fine-tunes LLMs for in-context operator learning. These methods typically involve custom architectures or task-specific training procedures tailored for PDE solving.
@@
Results
This section investigates the ability of state-of-the-art open-weight foundation LLMs to continue the spatiotemporal dynamics of PDEs. We focus our main analysis on the Allen–Cahn equation [29], a nonlinear PDE modeling phase separation in multi-component metal alloy systems. To assess generality, we additionally evaluate the Fisher–KPP equation [30, 31], another nonlinear PDE modeling population growth and diffusion, together with two representative linear PDEs: the heat equation (a parabolic diffusion model) and the wave equation (a hyperbolic wave propagation model) [19]. Across these families, with random initial conditions and varying boundary condition types, we observe qualitatively consistent ICL behavior. Remarkably, LLM rollouts also approximately conserve total thermal energy in the heat equation under Neumann boundaries, indicating that zero-shot ICL captures not only spatiotemporal dynamics but also structural invariants of PDEs. Full results and discussion of these additional PDE experiments are provided in Supplementary Material Sect. 3.
PDE Setup.
By coupling reaction and diffusion dynamics, the Allen–Cahn equation induces strong interactions across space and time, making it a challenging yet physically interpretable testbed for assessing whether LLMs capture genuinely spatiotemporal structure rather than merely extrapolating along a single dimension. The system is defined on the interval [1,1][-1,1] with Dirichlet boundary conditions u(1,t)=u(1,t)=1u(-1,t)=u(1,t)=-1 and random initial conditions $u(x,0)=u_{0}(x)$ (details of initial condition generation in Supplementary Material Sect. 1). Explicitly, the Allen–Cahn PDE is:
Here, $\partial_{t}$ and $\partial_{xx}$ denote the temporal and second-order spatial derivatives, respectively. Adopting standard parameter choices [32, 33], we set the diffusion coefficient to $\epsilon^{2}=0.001$ and use a double-well potential $f(u)=2(u^{3}-u)$ for the nonlinear reaction term. The solution is evaluated on a uniform spatiotemporal grid $\smash{\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}}}$, where $\smash{\{x_{i}\}_{i=1}^{N_{\mathrm{X}}}}$ are $N_{\mathrm{X}}$ evenly spaced interior points and $\smash{\{t_{j}\}_{j=0}^{N_{\mathrm{T}}}}$ are $N_{\mathrm{T}}+1$ evenly spaced time levels, with $T=0.5$.
Numerical Benchmarks.
To contextualize the predictive structure learned by LLMs, we compare their outputs to two classical finite difference methods: the forward time, centered space (FTCS) scheme, which is fully explicit, and an implicit-explicit (IMEX) scheme that treats diffusion implicitly and the nonlinear reaction term explicitly [34]. In contrast to developing new PDE solvers, our focus is on analyzing how pretrained foundation LLMs extrapolate PDE-based spatiotemporal dynamics in-context, without fine-tuning. Notably, their predictions can achieve surprising accuracy relative to standard numerical benchmarks, generalizing across varied initial conditions and discretizations. These results position PDEs as effective vehicles for examining the inductive biases and generalization behaviors of LLMs.
Analysis Task Overview.
We analyze how LLMs generalize in autoregressively continuing PDE dynamics through two analytical lenses: (i) a truncation-error perspective, motivated by local and global error analysis in numerical PDEs [35, 36], examining how prediction accuracy depends on discretization, rollout horizon, and model size; and (ii) a “systems-level” perspective, investigating how LLMs internalize and extrapolate PDE structure during ICL via entropy-based uncertainty measures. The One-Step Prediction section analyzes one-step prediction, where accuracy improves with longer temporal context but degrades with finer spatial discretization. Multi-Step Rollouts subsection analyzes multi-step rollouts, showing algebraic error growth with rollout horizon, analogous to global error accumulation in numerical solvers. The Uncertainty Evolution and Learning Stages section analyzes token-level uncertainty, revealing a consistent three-stage ICL progression: syntax mimicry, high-entropy exploration, and confident prediction. Prediction quality is evaluated using the Root Mean Square Error (RMSE), computed by aggregating errors over all spatial grid points $\{x_{i}\}_{i=1}^{N_{\mathrm{X}}}$ per time step $\{t_{j}\}_{j=1}^{N_{\mathrm{T}}}$:
RMSE measures overall predictive accuracy. Results using the Maximum Absolute Error, capturing worst-case deviation, show qualitatively similar trends (Supplementary Material Sect. 2). To ensure that our error metrics capture physically meaningful discrepancies—rather than differences between raw 3-digit quantized integer tokens—we compare each predicted solution $\hat{u}(x_{i},t_{j})$ against a floating-point reference solution $\tilde{u}(x_{i},t_{j})$. The reference $\tilde{u}(x_{i},t_{j})$ is computed on a highly refined finite-difference grid and passed through the same quantization–reconstruction pipeline (see Methods section) to ensure a consistent evaluation basis. LLM predictions, generated as token sequences, are likewise reconstructed into floating-point form before error evaluation. Since classical solvers operate on floating-point data, we apply the same quantization–reconstruction process to their initial conditions, ensuring that both LLM-based and classical methods operate on inputs of matched precision. Errors are then averaged over multiple random initial conditions, with further details in Supplementary Material Sect. 2.
Model Setup.
Our primary results focus on base pretrained models from the Meta Llama-3 family: Llama-3.1-8B, Llama-3.2-3B, and Llama-3.2-1B. We also evaluate their instruction-tuned counterparts optimized for dialogue (Supplementary Material Sect. 4), which display qualitatively similar trends. For broader comparison, we include Microsoft Phi-4 and Hugging Face SmolLM3, which exhibit similar behaviors with differences mainly in error magnitude, in Supplementary Material Sect. 5.
Figure 2: In-context error scaling with temporal discretization (left) and spatial discretization (right). The top axes show $N_{\mathrm{T}}$ and $N_{\mathrm{X}}$, while the bottom axes show the equivalent LLM context and output lengths $N_{\mathrm{Tokens}}$, respectively. RMSE decreases with longer context, converging in the extended-context regime, toward the local truncation behavior of first-order-in-time solvers (FTCS, IMEX). In contrast, errors grow with output length, following a capacity-dependent generalization trend. Shaded regions show 95% confidence intervals over 50 random initial conditions. The gray dotted line indicates the unavoidable quantization error floor defined in the Methods section.
One-Step Prediction
We evaluate one-step prediction error under the following setup: given a discretized PDE solution from the initial condition up to one step before the final time, $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}-1}$, the model predicts the terminal step $\{u(x_{i},t_{N_{\mathrm{T}}})\}_{i=1}^{N_{\mathrm{X}}}$ across all spatial grid points. This prediction target is employed in order to isolate the effects of temporal and spatial discretization on accuracy. To ensure the task remains non-trivial, we select discretizations where solutions vary significantly between time steps, avoiding degenerate cases in which the model could succeed through trivial pattern repetition. Supplementary Material Sect. 7 provides empirical validation of this design choice.
Longer Context Length Improves Prediction Accuracy.
We analyze how one-step prediction accuracy varies with input context length, measured by the number of observed time steps $N_{\mathrm{T}}$ provided to the LLM during ICL. Fixing the spatial discretization at $N_{\mathrm{X}}=14$, we vary $N_{\mathrm{T}}$ from 2 to 40, spanning minimal to moderately informative temporal contexts. As $N_{\mathrm{T}}$ increases, the LLM receives a longer input sequence, and RMSE decreases consistently (Fig. 2, left). On a log–log scale, the error curves exhibit approximate $\mathcal{O}(1/N_{\mathrm{T}})$ decay rate, closely resembling the convergence behavior of first-order-in-time solvers such as FTCS and IMEX. This suggests that LLMs exhibit inductive biases analogous to local truncation error in classical numerical methods.
Closer examination of Fig. 2 (left) reveals three distinct stages as the temporal context increases. The emergence and evolution of these stages—and their connection to prediction uncertainty—are further analyzed in the Uncertainty Evolution and Learning Stages section. In the limited-context stage ($N_{\mathrm{T}}=2$), LLMs exhibit substantially higher errors compared to classical solvers. This behavior arises from surface-level pattern imitation in the solution format, rather than learning the underlying dynamical structure, as examined in detail in the Uncertainty Evolution and Learning Stages section. In the moderate-context stage ($2<N_{\mathrm{T}}<10$), errors decay more rapidly than those of standard numerical benchmarks such as FTCS and IMEX, suggesting that LLMs move beyond surface-level pattern imitation and begin to internalize aspects of the governing PDE dynamics. Finally, in the extended-context stage ($N_{\mathrm{T}}\geq 10$), error decay closely matches that of classical first-order solvers, indicating that LLMs are effectively leveraging the spatiotemporal structure in a numerically grounded way. In this stage, Llama-3.1-8B consistently matches, or in some cases exceeds, the accuracy of classical solvers. Overall, this reveals an empirical in-context scaling law: increasing the input context length consistently improves prediction accuracy, reflecting the LLM’s increasing ability to internalize and extrapolate latent PDE dynamics at fixed spatial discretization.
Longer Output Length Degrades Prediction Accuracy.
We analyze how prediction accuracy varies with the number of spatial discretization points $N_{\mathrm{X}}$. Since LLMs predict the solution at all spatial grid points for a given time step, increasing $N_{\mathrm{X}}$ directly results in a proportionally longer output sequence. To isolate this effect, we vary $N_{\mathrm{X}}$ across 20 evenly spaced values from 2 to 40, while fixing the temporal context at $N_{\mathrm{T}}=50$. As shown in Fig. 2 (right), RMSE grows consistently with $N_{\mathrm{X}}$, following an approximate $\mathcal{O}(N_{\mathrm{X}})$ scaling law on a log–log scale. The error growth is steepest for the smaller Llama-3.2-1B model, while the larger Llama-3.1-8B shows slower growth, indicating improved robustness to output length within the Llama-3 family (see Supplementary Material Sect. 6 for architectural and size-related details of the Llama-3 models). This behavior stands in sharp contrast to classical finite-difference solvers, where increasing spatial resolution typically does not raise error under a stable scheme. For LLMs, however, outputs are generated as flat autoregressive token sequences: larger $N_{\mathrm{X}}$ leads to longer, more complex outputs that must be generated without access to the underlying PDE, placing growing demands on the model’s ICL capacity.
These findings reveal a second empirical in-context scaling law: finer spatial discretization produces longer outputs and degrades prediction accuracy under fixed input context. The effect scales strongly with model size within the Llama-3 family. Smaller models face more pronounced performance drops, indicating that limited ICL capacity constrains generalization at finer spatial discretizations.
Multi-Step Rollouts
We examine LLMs’ capacity to continue PDE solutions over multiple time steps based solely on in-context input. For a rollout of $N_{\mathrm{T}}$ time steps, we partition the serialized sequence into a context segment of $\lfloor\frac{2}{3}N_{\mathrm{T}}\rfloor$ steps and a prediction segment with the remainder. For $N_{\mathrm{T}}=25$, this yields 16 context steps (including the initial condition), $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},15}$, which the LLM uses to autoregressively generate $10$ prediction steps, $\{u(x_{i},t_{j})\}_{i=1,j=16}^{N_{\mathrm{X}},25}$, without access to intermediate ground truth. This 2:1 context-to-prediction ratio strikes a balance between providing sufficient context and posing a nontrivial extrapolation challenge. We assess model behavior via (i) representative rollouts from single random initial conditions, and (ii) average error trends over random initial conditions to quantify how prediction error accumulates over the prediction horizon.
Figure 3: Multi-step prediction for randomly sampled initial conditions of two PDEs. The first row shows the Allen–Cahn equation, and the second shows the wave equation ($c=0.3$; see Supplementary Material Sect. 3). In each case, to the left of the dashed line corresponds to the input context provided to the LLM, and to the right corresponds to a 10-step autoregressive continuation from a single generation of each model. Classical finite difference solvers (FTCS for Allen–Cahn, leapfrog for wave) solve the corresponding initial value problem using the final in-context time slice as the initial condition, and advance the solution for 10 steps using the same spatial and temporal discretization as the LLMs. The final three columns report pointwise absolute errors relative to the reference solution.
Qualitative Multi-Step Rollouts.
Figure 3 illustrates representative multi-step rollouts for the Allen–Cahn and wave equations, showing that LLMs can sustain coherent, qualitatively accurate predictions over a 10-step horizon. This is notable because the foundation models are not specialized PDE solvers and lack training-time exposure to the discretized PDE solutions; initial conditions are randomly sampled at inference time. The Llama-3.1-8B model closely tracks the evolution, capturing nonlinear reaction–diffusion dynamics and finite-speed wave propagation without collapsing to trivial behavior or diverging. When prompted with sufficiently long input context, the model can approximate and extrapolate PDE dynamics purely through autoregressive token-level inference. In contrast, Llama-3.2-1B shows larger deviations and fails to preserve spatiotemporal structure over extended rollouts (see Supplementary Material Sect. 9, for analysis of error patterns and capacity limits). Visualizations for Llama-3.2-3B, additional numerical benchmarks, and further random initial conditions are provided in Supplementary Material Sect. 8.
Quantitative Multi-Step Error Growth.
To characterize error growth over the prediction horizon, we repeat the multi-step rollout procedure for the Allen–Cahn equation across 20 randomly sampled initial conditions. Averaging over initializations reveals a consistent algebraic increase in RMSE with rollout length, as shown on a log–log scale in Fig. 4. This resembles global error accumulation in classical finite-difference solvers operating under stable discretizations, where local truncation errors compound in a controlled manner over time. Crucially, none of the models exhibit divergent or unstable behavior across the 10-step horizon; error growth remains bounded. Moreover, a comparison with naive autoregressive models further shows that the extrapolation behavior exhibited by LLMs cannot be attributed to simple continuation methods (see Fig. S17 in the Supplementary Material). These findings underscore the capacity of LLMs to continue PDE dynamics via ICL across diverse initial conditions, sustaining coherent predictions over extended horizons—a fundamentally nontrivial task given only in-context information, without prompting or access to governing equations. Similar error-growth trends are observed for the wave equation and other PDEs, as detailed in Supplementary Material Sect. 3.
Uncertainty Evolution and Learning Stages
We now move beyond truncation-like error analysis to examine how LLMs internalize PDE dynamics through ICL. We focus on predictive uncertainty and generation behavior in the one-step prediction task introduced in the One-Step Prediction section. Since the multi-step rollouts recursively apply the one-step prediction process, we defer a parallel analysis of uncertainty accumulation in that setting to Supplementary Material Sect. 10. We vary context length via temporal discretization ($N_{\mathrm{T}}$) and output length via spatial discretization ($N_{\mathrm{X}}$), and analyze how these factors shape the model’s token-level uncertainty. Given input $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}-1}$, the model predicts $\{u(x_{i},t_{N_{\mathrm{T}}})\}_{i=1}^{N_{\mathrm{X}}}$. To quantify predictive uncertainty, we compute the Shannon entropy [37] of the model’s softmax distribution at each spatial value token and average across space:
where $\mathcal{V}$ denotes the tokenizer’s vocabulary, and $p(y\mid x_{i},N_{\mathrm{T}})$ is the predicted probability of token $y$ at spatial location $x_{i}$, given $N_{\mathrm{T}}$ prior time steps in the serialized input. See Supplementary Material Sect. 2 for implementation details.
Figure 4: Multi-step rollout error trends. RMSE grows algebraically with prediction steps $n$ (top axis) and equivalent LLM output length $N_{\mathrm{Tokens}}$ (bottom axis). Left: rollout from a single random initial condition (as in Fig. 3). Right: average over 20 random initial conditions. Error growth rates are estimated via log–log fits and reported on the right. Shaded regions denote 95% confidence intervals (left: across 20 repeated LLM runs; right: across 20 initial conditions).
Emergent Learning Stages Revealed by Entropy Evolution.
Figure 5a shows the evolution of mean spatial entropy, $\bar{H}$, as a function of $N_{\mathrm{T}}$. A distinct rise-and-fall pattern reveals three emergent stages of ICL: an initial syntax-dominated stage, a transitional exploratory stage, and a final stage of consolidation and refinement:
  1. 1)
    Syntax-Only (limited context, e.g., $N_{\mathrm{T}}=2$; Fig. 5c, first row): While mean spatial entropy $\bar{H}$ exhibits some variation across model sizes, prediction error remains consistently high. Separator tokens (e.g., commas) are predicted with near-perfect confidence (see Fig. 5d). In contrast, spatial value tokens act as generic placeholders, with little correspondence to underlying PDE dynamics, yielding deterministic yet physically implausible predictions. This indicates that syntax is acquired before any meaningful understanding of the PDE dynamics emerges.
  2. 2)
    Exploratory (moderate context, e.g., $2<N_{\mathrm{T}}<10$; Fig. 5c, second row): Entropy reaches its peak across model sizes, indicating increased uncertainty and broader spatial token distributions. Meanwhile, prediction accuracy improves rapidly, and outputs begin to align with true PDE dynamics. This stage marks a transition from merely capturing surface-level syntax to beginning to internalize spatiotemporal dynamics.
  3. 3)
    Consolidation (extended context, e.g., $N_{\mathrm{T}}\geq 10$; Fig. 5c, third row): As context length increases further, $\bar{H}$ decreases, reflecting sharper and more confident spatial token distributions. Prediction accuracy continues to improve, though with less profound gains compared to the exploratory stage. The model’s predictions increasingly reflect coherent and physically meaningful PDE dynamics.
These stages reveal a consistent ICL progression: 1) syntax acquisition, 2) exploratory numerical behavior, and 3) convergence to accurate predictions. This progression suggests that LLMs develop structured internal representations of PDE dynamics purely through in-context exposure, without explicit access to governing equations or language prompting.
Uncertainty Growth with Output Length.
While the previous analysis focused on how input context length affects ICL and predictive uncertainty, we now examine how spatial discretization ($N_{\mathrm{X}}$) affects prediction confidence. Fixing the temporal context at $N_{\mathrm{T}}=50$, we vary $N_{\mathrm{X}}$ and compute the mean spatial entropy $\bar{H}$. As shown in Fig. 5b, $\bar{H}$ increases steadily with larger $N_{\mathrm{X}}$, reflecting growing uncertainty for longer spatial outputs. This trend mirrors the error scaling in Fig. 2, where RMSE increases with $N_{\mathrm{X}}$ under fixed input context. Notably, the smaller Llama-3.2-1B exhibits the steepest entropy growth, while the larger Llama-3.1-8B shows the slowest, indicating greater robustness to output length. These findings reveal a close empirical link between model uncertainty and prediction error: longer output sequences from finer spatial discretization lead to both higher entropy and reduced predictive accuracy. Within the Llama-3 family, larger models consistently maintain higher confidence and accuracy over extended output lengths. This suggests that model size plays a critical role in enabling accurate extrapolation of learned PDE dynamics across finer spatial discretizations—a capacity that smaller models fail to maintain.
Figure 5: Three-stage ICL progression and the evolution of predictive uncertainty. (a–b) Mean spatial entropy $\bar{H}$ vs. (a) temporal context length $N_{\mathrm{T}}$ at fixed $N_{\mathrm{X}}=14$ and (b) output length $N_{\mathrm{X}}$ at fixed $N_{\mathrm{T}}=50$. Shaded regions: 95% confidence intervals over 50 random initial conditions. (c) Token-level softmax distributions at three ICL stages: syntax-only ($N_{\mathrm{T}}=2$), exploratory ($N_{\mathrm{T}}=5$), and consolidation ($N_{\mathrm{T}}=20$), extracted from Llama-3.1-8B for the same initial condition as the multi-step rollout example. Top 8 tokens (by probability) are shown per spatial position; only odd positions are displayed, with full results in Supplementary Material Sect. 11. (d) Softmax over separator tokens. Early, high-confidence delimiter predictions are the signature of the syntax-only stage: the model acquires and stabilizes delimiter syntax with minimal context, and this high-confidence behavior over separators is preserved as the model subsequently learns the PDE dynamics.
Discussion
We show that text-trained LLMs can extrapolate PDE dynamics in-context in a zero-shot setting, without any fine-tuning or natural language prompting. Their performance exhibits clear in-context scaling laws: accuracy improves with longer temporal context, degrades with finer spatial discretization, and error grows algebraically under multi-step rollouts. Entropy analysis further reveals a three-phase progression—from syntax imitation, to exploratory uncertainty, to stabilized predictions—highlighting emergent mechanisms underlying ICL. Together, these findings suggest that LLMs can internalize nontrivial aspects of PDE dynamics purely from in-context data, demonstrating emergent generalization capabilities in zero-shot inference.
Our study focuses on time-dependent PDEs with real-valued solutions under full observation. Extending this framework to stationary PDEs (e.g., Poisson), complex-valued systems (e.g., Schrödinger), and partially observed or noisy dynamics could reveal complementary behaviors. Another direction is to investigate how LLMs develop or express numerical priors when applied to time-dependent PDEs in higher spatial dimensions, and whether new ICL behaviors emerge as spatial complexity increases, potentially offering additional insight into how LLMs represent space and time[38]. While our experiments deliberately avoid natural-language prompting or symbolic information about the governing PDE, boundary conditions, or initial conditions in order to isolate intrinsic zero-shot capability, incorporating prior knowledge such as physics-aware qualitative or symbolic descriptions [39, 12] may serve as an informative cue, revealing how explicit structure shapes in-context reasoning and inductive biases. Beyond these extensions, a key challenge is to characterize the internal representations and compositional structures that support generalization over spatiotemporal dynamics in autoregressive token space.
Methods